# Predicting the rate of fast radio bursts in globular clusters from binary black hole observations

### v1

@@ 1 Introduction
Fast radio bursts (FRBs) are bright, millisecond-long radio flares that have been observed to originate from extragalactic sources, with some sources giving rise to repeating bursts (Lorimer et al., 2007; Spitler et al., 2014a; Marcote et al., 2017). In recent years, the discovery rate of these mysterious signals has accelerated. For instance, the Parkes radio telescope is credited to numerous FRB detections including the first detection (namely the Lorimer burst) in 2007 (Lorimer et al., 2007; Keane et al., 2012; Thornton et al., 2013; Burke-Spolaor & Bannister, 2014; Petroff et al., 2015; Ravi et al., 2015; Champion et al., 2016). Detections have also been made by the Arecibo telescope (Spitler et al., 2014b) and the Green Bank telescope (Masui et al., 2015). The Canadian Hydrogen Intensity and Mapping Experiment (CHIME) Telescope discovered and catalogued more than 500 FRBs in 2021 (CHIME/FRB Collaboration et al., 2021). The Australian Square Kilometre Array Pathfinder (ASKAP) recently reported the detection of a very distant FRB at redshift $z>1$ (Ryder et al., 2023). Despite the numerous detections, there is no common consensus amongst the community on the sources of these bursts.
The redshift evolution of FRB rates may provide strong indications of their origins. On one hand, some studies argue that the rate of FRBs closely follows the star formation history (SFH; e.g., James et al., 2022; Bhardwaj et al., 2023; Wang & van Leeuwen, 2024), implying a connection to short-lived, massive stars. This is supported by the recent detection of FRB200428 from the Galactic soft gamma repeater SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020), which clearly shows that some FRBs are generated by violent activities on the surface of young magnetars (Thompson & Duncan, 1995; Kashiyama et al., 2013; Popov & Postnov, 2013; see Petroff et al., 2022, for a review). Magnetars are young neutron stars (NS) with extremely strong magnetic fields ($\sim 10^{14}$ G), traditionally thought to form by the collapse of massive stars in core-collapse supernovae. Magnetars radiate their energy over a span of a few thousand years, which causes their magnetic fields to decay until they reache regular NS levels (see Kaspi & Beloborodov, 2017, for a review of magnetars). Hence, if magnetars formed via core-collapse supernovae contribute to the majority of the observed FRBs, the FRB rate as a function of redshift must trace the SFH of the Universe, which increases with increasing redshift in the local Universe before peaking at $z\approx 2$.
@@ 2 Population of Globular Clusters
We utilize a large grid of realistic simulations of GCs run with the state-of-the-art Cluster Monte Carlo (CMC) $N$-body code (Kremer et al., 2020b) to study the BWD coalescence rates from GCs. We assume the GC population follows the constraints derived from GW observations (Fishbach & Fragione, 2023).
@@ 2.1 Simulations of globular clusters
This gives a total of $144$ simulations with every combination of the above-mentioned properties. Tracked in each simulation are the binary and dynamical interactions of the stellar constituents. The stellar initial mass function is assumed to follow the Kroupa standard broken power law (Kroupa, 2001) between $0.08\ M_{\astrosun}$ and $150\ M_{\astrosun}$. The initial binary fraction is set to 5%.
@@ 3.1 Predicting the rate of binary white dwarf coalescences
Our approach was verified on the BBH coalescences data from the CMC simulations. For the default values of $S_{\text{param}}$ mentioned in Section 2.2, we obtain a BBH coalescence rate curve that peaks at $z=1.4$. When normalised to $10\ \text{Gpc}^{-1}\ \text{yr}^{-1}$ at $z=0$, the peak rate equals $20\ \text{Gpc}^{-1}\text{yr}^{-1}$. This is in close agreement with the results obtained by Ye & Fishbach (2024) on the same BBH coalescence data (the black curve in Figure 2 of their paper). Running similar tests with GC parameter values used by Fishbach & Fragione (2023) we were able to re-obtain all the BBH coalescence rates given in Figure 2 of their paper.
@@ 3.2 Rate of fast radio bursts
Figure 3: FRB formation/source rates as a function of redshift. SFH Models: Bright pink spaced double dot dashed: SFH model from Shin et al. (2023). Orange: SFH model from Zhang & Zhang (2022). Purple dense double dot dashed: Accumulated model from Zhang & Zhang (2022). Time Delayed Models: Red dashed: Lognormal delay model with a central value of 10 Gyr from Zhang & Zhang (2022). Blue dotted: Best-fit model to CHIME data from Zhang et al. (2024). Green solid: Hybrid model that is $20\%$ SFH and $80\%$ lognormal delay with a mean of 13 Gyr from Zhang & Zhang (2022). Grey star scatter: Best-fit model of CHIME data from Hashimoto et al. (2022). Teal plus scatter: Best-fit model of CHIME data from Chen et al. (2024). Here the lognormal delay model, hybrid model, accumulated model and SFH model from Zhang & Zhang (2022) were scaled up using a multiplicative factor of $1.25\times 10^{4}\ {\rm Gpc^{-3}\,yr^{-1}}$, which equals the value of the FRB rate from Zhang et al. (2024) at redshift $z=0$.
@@ 3.3 Comparing binary white dwarf coalescence rates with fast radio burst rates
Figure 4: Comparisons between the inferred BWD coalescence rates from GCs outlined in Section 3.1 and the FRB formation/source rates as a function of redshift. In all the panels, the upper and lower thick solid lines represent the $95^{\text{th}}$ percentile curve and the $5^{\text{th}}$ percentile curve for the BWD coalescence rates, respectively. The solid black curve shows the mean BWD coalescence rate in each panel, and the dashed grey line shows the BWD coalescence rate obtained using the default GC parameters which correspond to a BBH coalescence rate of $10\ \text{Gpc}^{-3}\ \text{yr}^{-1}$ at $z=0$ (see Section 2.2). The legends for the FRB rate curves are as follows. Red dashed: Lognormal delay model with a central value of 10 Gyr from Zhang et al. (2021). Blue dotted: Best-fit model to CHIME data from Zhang et al. (2024). Green solid: Hybrid model where there is 20% SFH and 80% lognormal delay with a mean of 13 Gyr from Zhang et al. (2021). Grey star scatter: Best-fit model for the CHIME data from Hashimoto et al. (2022). Teal plus scatter: Best-fit model of the CHIME data from Chen et al. (2024).The cyan star is the rate of BWD coalescences at $z=0$ as determined by Kremer et al. (2023b).
@@ 4.2 Uncertainties in binary white dwarf coalescences
Similarly, BWD coalescences in galactic fields may also produce FRB-emitting NSs. The coalescence rate of two super-Chandrasekhar CO WDs in the field is roughly an order of magnitude lower than the rates of Type Ia supernovae ($\sim 10^{4}\,{\rm Gpc^{-1}\,yr^{-1}}$; Badenes & Maoz, 2012; Maoz & Mannucci, 2012, e.g.,), which is close to the FRB source rates in the local Universe. However, Ruiter et al. (2009) showed that the delay time of BWD coalescences in the field is about 1 Gyr, similar to the delay time of the cluster BWD mergers from primordial binaries ($\sim 100~{}$Myr). Therefore, the redshift evolution of field BWD coalescences likely follows the SFH more closely than old stellar populations. This suggests that if BWD coalescences from binary evolution are indeed one of the dominant mechanism for producing FRB sources both in the field and in clusters, the FRB formation rates would likely peak at higher redshift rather than decreasing with increasing redshift.
@@ 5 Conclusions
We are grateful to Kyle Kremer and Mohit Bhardwaj for useful discussions. C.S.Y. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) DIS-2022-568580. M.F. acknowledges support from NSERC RGPIN-2023-05511, the University of Toronto Connaught Fund, and the Alfred P. Sloan Foundation.

### v2

@@ 1 Introduction
Fast radio bursts (FRBs) are bright, millisecond-long radio flares that have been observed to originate from extra-Galactic sources, with some sources giving rise to repeating bursts (Lorimer et al., 2007; Spitler et al., 2014; Marcote et al., 2017). In recent years, the discovery rate of these mysterious signals has accelerated. For instance, the Parkes radio telescope is credited to numerous FRB detections including the first detection (namely the Lorimer burst) in 2007 (Lorimer et al., 2007; Keane et al., 2012; Thornton et al., 2013; Burke-Spolaor & Bannister, 2014; Petroff et al., 2015; Ravi et al., 2015; Champion et al., 2016). Detections have also been made by the Arecibo telescope (Spitler et al., 2014) and the Green Bank telescope (Masui et al., 2015). The Canadian Hydrogen Intensity and Mapping Experiment (CHIME) Telescope discovered and catalogued more than 500 FRBs in 2021 (CHIME/FRB Collaboration et al., 2021). The Australian Square Kilometre Array Pathfinder (ASKAP) recently reported the detection of a very distant FRB at redshift $z>1$ (Ryder et al., 2023). Despite the numerous detections, there is no common consensus amongst the community on the sources of these bursts.
The redshift evolution of FRB rates may provide strong indications of their origins. On one hand, some studies argue that the rate of FRBs closely follows the star formation history (SFH; e.g., James et al., 2022; Bhardwaj et al., 2024b; Wang & van Leeuwen, 2024), implying a connection to short-lived, massive stars. This is supported by the recent detection of FRB200428 from the Galactic soft gamma repeater SGR 1935+2154 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020), which clearly shows that some FRBs are generated by violent activities on the surface of young magnetars (Thompson & Duncan, 1995; Kashiyama et al., 2013; Popov & Postnov, 2013; see Petroff et al., 2022, for a review). Magnetars are young neutron stars (NS) with extremely strong magnetic fields ($\sim 10^{14}$ G), traditionally thought to form by the collapse of massive stars in core-collapse supernovae. Magnetars radiate their energy over a span of a few thousand years, which causes their magnetic fields to decay until they reach regular NS levels (see Kaspi & Beloborodov, 2017, for a review of magnetars). Hence, if magnetars formed via core-collapse supernovae contribute to the majority of the observed FRBs, the FRB rate as a function of redshift must trace the SFH of the Universe, which increases with increasing redshift in the local Universe before peaking at $z\approx 2$.
@@ 2 Population of Globular Clusters
We utilize a large grid of realistic simulations of GCs run with the state-of-the-art Cluster Monte Carlo (CMC) $N$-body code (Kremer et al., 2020b) to study the BWD coalescence rates from GCs. We assume the GC population follows the constraints derived from GW observations (Fishbach & Fragione, 2023, their Figure B1 and B2). These constraints apply to a set of parameters that define the GC population properties, which are further described in Section 2.2 below.
@@ 2.1 Simulations of globular clusters
This gives a total of $144$ simulations with every combination of the above-mentioned properties. Tracked in each simulation are the binary and dynamical interactions of the stellar constituents. The stellar initial mass function is assumed to follow the Kroupa standard broken power law (Kroupa, 2001) between $0.08\ M_{\astrosun}$ and $150\ M_{\astrosun}$. The initial binary fraction for all stars is set to $5\%$ and the mass of the secondary star is drawn from a flat distribution in mass ratio in the range $[0.1,1]$ (e.g., Duquennoy & Mayor, 1991).
@@ 3.1 Predicting the rate of binary white dwarf coalescences
Our approach was verified on the BBH coalescences data from the CMC simulations. For the default values of $S_{\text{param}}$ mentioned in Section 2.2, we obtain a BBH coalescence rate curve that peaks at $z=1.4$. When normalised to $10\ \text{Gpc}^{-3}\ \text{yr}^{-1}$ at $z=0$, the peak rate equals $20\ \text{Gpc}^{-3}\text{yr}^{-1}$. This is in close agreement with the results obtained by Ye & Fishbach (2024) on the same BBH coalescence data (the black curve in Figure 2 of their paper). Running similar tests with GC parameter values used by Fishbach & Fragione (2023) we were able to re-obtain all the BBH coalescence rates given in Figure 2 of their paper.
@@ 3.2 Rate of fast radio bursts
Figure 3: FRB formation/source rates as a function of redshift. SFH Models: A: Accumulated model from Zhang & Zhang (2022). B: SFH model from Zhang & Zhang (2022). C: SFH model from Shin et al. (2023). Time Delayed Models: D: Best-fit model of CHIME data from Hashimoto et al. (2022). E: Best-fit model of CHIME data from Chen et al. (2024). F: Hybrid model that is $20\%$ SFH and $80\%$ lognormal delay with a mean of 13 Gyr from Zhang & Zhang (2022). G: Lognormal delay model with a central value of 10 Gyr from Zhang & Zhang (2022). H: Best-fit model to CHIME data from Zhang et al. (2024). Here the lognormal delay model, hybrid model, accumulated model and SFH model from Zhang & Zhang (2022) were scaled up using a multiplicative factor of $1.25\times 10^{4}\ {\rm Gpc^{-3}\,yr^{-1}}$, which equals the value of the FRB rate from Zhang et al. (2024) at redshift $z=0$.
@@ 3.3 Comparing binary white dwarf coalescence rates with fast radio burst rates
Figure 4: Comparisons between the inferred BWD coalescence rates from GCs outlined in Section 3.1 and the FRB formation/source rates as a function of redshift. In all the panels, the upper and lower thick solid lines represent the $95^{\text{th}}$ percentile curve and the $5^{\text{th}}$ percentile curve for the BWD coalescence rates, respectively. The solid black curve shows the mean BWD coalescence rate in each panel, and the dashed grey line shows the BWD coalescence rate obtained using the default GC parameters which correspond to a BBH coalescence rate of $10\ \text{Gpc}^{-3}\ \text{yr}^{-1}$ at $z=0$ (see Section 2.2). The legends for the FRB rate curves are as follows. D: Best-fit model of CHIME data from Hashimoto et al. (2022). E: Best-fit model of CHIME data from Chen et al. (2024). F: Hybrid model that is $20\%$ SFH and $80\%$ lognormal delay with a mean of 13 Gyr from Zhang & Zhang (2022). G: Lognormal delay model with a central value of 10 Gyr from Zhang & Zhang (2022). H: Best-fit model to CHIME data from Zhang et al. (2024). The cyan star is the rate of BWD coalescences at $z=0$ as determined by Kremer et al. (2023b).
@@ 4.2 Uncertainties in binary white dwarf coalescences
Similarly, BWD coalescences in galactic fields may also produce FRB-emitting NSs. The coalescence rate of two super-Chandrasekhar CO WDs in the field is roughly an order of magnitude lower than the rates of Type Ia supernovae ($\sim 10^{4}\,{\rm Gpc^{-3}\,yr^{-1}}$; Badenes & Maoz, 2012; Maoz & Mannucci, 2012, e.g.,), which is close to the FRB source rates in the local Universe. However, Ruiter et al. (2009) showed that the delay time of BWD coalescences in the field is about 1 Gyr, similar to the delay time of the cluster BWD mergers from primordial binaries ($\sim 100~{}$Myr). Therefore, the redshift evolution of field BWD coalescences likely follows the SFH more closely than old stellar populations. This suggests that if BWD coalescences from binary evolution are indeed one of the dominant mechanism for producing FRB sources both in the field and in clusters, the FRB formation rates would likely peak at higher redshift rather than decreasing with increasing redshift.
While we have fixed the initial binary fraction to 5% in our cluster simulations, we do not expect the results to be affected significantly by this choice. The number of BBH mergers is mostly dominated by the dynamically-produced systems (e.g., Chatterjee et al., 2017). A larger initial binary fraction likely also has negligible effects on the number of BWD mergers (see, e.g., Kremer et al., 2021b, their Section 7.3). Increasing the initial binary fraction from $5\%$ to $50\%$ and varying the initial semi-major axis distribution, for example, will likely increase the number of BWD mergers by a factor of 10 at most, which does not affect the conclusions of this study.
@@ 5 Conclusions
We are grateful to Kyle Kremer and Mohit Bhardwaj for useful discussions and to the anonymous referee for helpful comments. C.S.Y. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) DIS-2022-568580. M.F. acknowledges support from NSERC RGPIN-2023-05511, the University of Toronto Connaught Fund, and the Alfred P. Sloan Foundation.